Two consecutive odd numbers total 32. Find the smaller.
Found 2 solutions by checkley77, katelynsaxe:
Answer by checkley77(12569)

Let x & x=2 be the 2 odd numbers.
x+x+2=32
2x=32-2
2x=30
x=30/2
x=15 ans. for the smaller number.
Proof:
15+15+2=32
32=32

Answer by katelynsaxe(2)

15,17 so 15 would be the smaller of the two. The answer is 15